# Inverse Constitutional AI: Compressing Preferences into Principles

### v1

@@ 6 Limitations and discussion
It is important to consider the limitations of our approach when interpreting our results. Firstly, we do not show causality—our generated principles correlate LLM annotators with the original annotations, but we cannot validate if these principles were used by the original annotators. Multiple constitutions may explain the data equally well (as in the Rashomon effect [Breiman, 2001]). Nonetheless, an undesirable principle correlating with annotations is concerning, even if the principle was not intentionally used. For example, in the ‘aligned’ variant of the AlpacaEval dataset, some generated constitutions include principles to prefer verbose or redundant responses (see Section C.2.1). While this principle was likely not consciously followed by the original annotators, its high support in the dataset may warrant further investigation and possible data cleaning. Secondly, constitutions represent a lossy compression—A constitution of a few principles is a simplification of the decision-making process underlying annotations. Some annotations may not be possible to reconstruct based on a simple constitution. [^9] Similarly, a country’s legal constitution is an ambiguous set of principles interpreted by judges, or LLMs in our case. Finally, preferences closely aligned to LLMs are challenging to test. If an LLM annotator is already highly aligned with the dataset annotations, improving its performance with a constitution is challenging. The constitutional reconstruction loss is most useful for evaluating principles orthogonal to or against the popular opinions internalized by the LLM. On already well-aligned models, the constitution may not improve performance, but it can still provide insights into the underlying preferences.
We hope our work will have a positive societal impact by helping better understand preference data, already widely used for fine-tuning and evaluation of popular LLMs. We emphasize that our generated constitutions cannot claim to re-construct an individual’s true reasoning process. The constitutions should be interpreted cautiously when working with human annotators to avoid potential negative implications. Similarly, the use of our approach for personalised LLMs should also be considered carefully. Overall, we believe the potential for positive impacts outweighs possible negative impacts.
Future work should focus on addressing these limitations, extending the capabilities of our approach, possibly using multi-modal models, and exploring new applications.
@@ 7 Conclusion
We have presented our work on the Inverse Constitutional AI (ICAI) problem: first defining the ICAI problem compressing preference data into a short list of natural language principles (or constitution). We then introduced an initial ICAI algorithm as a first approach to generate such constitutions. We demonstrated the effectiveness of our approach in experiments across three different types of datasets: (a) synthetic data to provide a proof of concept; (b) AlpacaEval data to show the applicability to compress human-annotated data and the possibility of transferring constitutions across model families; and (c) Chatbot Arena data to illustrate the generation of personal constitutions. We hope that our approach can improve both our understanding and the usefulness of widely-used feedback data. Potential use-cases of our interpretable and editable constitutions include: highlighting issues with datasets, creating interpretable alternatives to black-box reward models, scaling human-annotated evaluation to new models and use-cases, and improving model customization via personal constitutions. We are excited for future work to explore these use-cases in more detail.
Acknowledgments and Disclosure of Funding
@@
We thank Benjamin Minixhofer, Max Bartolo, Tom Hosking, Hritik Bansal, the RECOG-AI team at the Leverhulme Centre for the Future of Intelligence (especially José Hernández-Orallo and Lorenzo Pacchiardi), Bill Marino and Jason Brown for their valuable feedback and discussions on early versions of this work. We also thank Anthropic for providing free research access to their models. Arduin Findeis was supported by a University of Cambridge School of Physical Sciences Award and by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (reference EP/S022961/1). This publication was further supported by LMUexcellent, funded by the Federal Ministry of Education and Research (BMBF) and the Free State of Bavaria under the Excellence Strategy of the Federal Government and the Länder as well as by the Hightech Agenda Bavaria.

### v2

@@ 6 Limitations
When interpreting our results, there are several limitations to our approach important to consider. Firstly, we do not show causality — our generated principles correlate LLM annotators with the original annotations, but we cannot validate if these principles were used by the original annotators. Multiple constitutions may explain the data equally well (as in the Rashomon effect (Breiman, 2001)). Nonetheless, an undesirable principle correlating with annotations is concerning, even if the principle was not intentionally used. For example, in the “aligned” variant of the AlpacaEval dataset, some generated constitutions include principles to prefer verbose or redundant responses (see Section H.2.1). While this principle was likely not consciously followed by the original annotators, its high support in the dataset may warrant further investigation and possible data cleaning. We further discuss this limitation in Appendix C. Secondly, constitutions represent a lossy compression — a constitution of a few principles is a simplification of the decision-making process underlying annotations. Some annotations may not be possible to reconstruct based on a simple constitution. This trade-off highlights the tension between interpretability and accuracy: concise, human-readable principles versus more complex representations. While ICAI could be adapted for richer constitutions to balance this trade-off, a black-box reward model may be preferable when maximizing accuracy is critical. Finally, preferences closely aligned to LLMs are challenging to test. If an LLM annotator is already highly aligned with the dataset annotations, improving its performance with a constitution is challenging. The constitutional reconstruction loss is most useful for evaluating constitutions orthogonal to or against the popular opinions internalized by the LLM. On already well-aligned models, the constitution may not improve performance, but it can still provide insights into the underlying preferences. Future work should focus on addressing these limitations, extending the capabilities of our approach, possibly using multi-modal models, and exploring new applications.
@@ 7 Conclusion
We have presented our work on the Inverse Constitutional AI (ICAI) problem: first defining the ICAI problem compressing preference data into a short list of natural language principles (or constitution). We then introduced an initial ICAI algorithm as a first approach to generate such constitutions. We demonstrated the effectiveness of our approach in experiments across four different types of datasets: (a) synthetic data to provide a proof-of-concept; (b) AlpacaEval data to show the applicability to compress human-annotated data and the possibility of transferring constitutions across model families; (c) Chatbot Arena data to illustrate the generation of personal constitutions; and (d) PRISM data to demonstrate the ability to provide possible explanations for previously observed group preferences. We hope that our approach can improve both our understanding and the usefulness of widely-used feedback data. Potential use cases of our interpretable and editable constitutions and corresponding meta data include: highlighting issues with datasets, creating interpretable alternatives to black-box reward models, scaling human-annotated evaluation to new models and use cases, and improving model customization via personal or group constitutions. We look forward to future research that explores these use cases in detail.
Ethics statement
We hope our work will have a positive societal impact by helping better understand preference data, already widely used for fine-tuning and evaluation of popular LLMs. We emphasize that our generated constitutions cannot claim to reconstruct an individual’s true reasoning process. Similar to other interpretability methods, an ICAI constitution’s principles may correlate with an individual’s annotations but no causal relationship between the constitution and the annotator’s reasoning can be proven. Further, our constitutions represent a notable compression of annotation considerations, which makes them highly interpretable but also means that they cannot reflect more multi-faceted decision making processes.
Thus, constitutions should be interpreted cautiously when working with human annotators to avoid potential negative implications. This is especially important when attempting to explain demographic preferences, as multiple possible explanations may correlate with the data and malicious actors could cherry-pick specific ones to make discriminatory statements or reinforce prior beliefs. Similarly, the use of our approach for personalized LLMs should also be considered carefully.
In general, we emphasize that our method can only provide information about specific preference annotation datasets rather than annotators’ reasoning processes more broadly. To mitigate the potential for misinterpretation of results, we include a corresponding warning in our algorithm implementation that is shown to the user whenever a constitution is generated with ICAI.
When using ICAI for certain downstream use-cases, such as annotation scaling for training and evaluation or generating personal constitutions, there exists a risk of harmful bias amplification. If harmful biases exist in the original preference dataset, the ICAI constitution may pick up on these and propagate them downstream. This risk of amplifying biases is counterbalanced, however, by the ability to edit and inspect the generated principles. This ability potentially helps avoid amplification of unintended biases when using ICAI in downstream applications. This potential visibility of biases in ICAI distinguishes our method from other widely-used methods using preference data, such as black-box reward models or aggregate evaluation statistics, which make such biases more difficult to detect. We recommend users of our method to always take a close look if generated constitutions are aligned with their own values and contain any potentially harmful biases before proceeding with downstream use-cases. Overall, we believe the potential for positive impacts outweighs possible negative impacts.
Reproducibility statement
We make the source code for our method and experiments publicly available at https://github.com/rdnfn/icai. Further, we attempted to add as many details in the paper as possible, including all prompts in Appendix G as well as a description of our synthetic data generation approach in Appendix I. For direct comparability, we also make numerical results available in Section F.9.
Acknowledgements
@@
We thank Benjamin Minixhofer, Max Bartolo, Tom Hosking, Hritik Bansal, the RECOG-AI team at the Leverhulme Centre for the Future of Intelligence (especially José Hernández-Orallo and Lorenzo Pacchiardi), Bill Marino and Jason Brown for their valuable feedback and discussions on early versions of this work. Further, we thank all reviewers for taking the time to read our work and to provide helpful feedback. We also thank Anthropic for providing free research access to their models. Arduin Findeis was supported by a University of Cambridge School of Physical Sciences Award and by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (reference EP/S022961/1). This publication was further supported by LMUexcellent, funded by the Federal Ministry of Education and Research (BMBF) and the Free State of Bavaria under the Excellence Strategy of the Federal Government and the Länder as well as by the Hightech Agenda Bavaria.